The accounting firm of Crawford and Associates has five senior partners. Yesterday the senior partners saw six, five, six, five, and nine clients, respectively.

a. Compute the mean and median number of clients seen by the partners. (Round your answers to 2 decimal places.)

Mean
Median

b. The data represents a _______ (Sample, population, grouped data, geometric mean)
Answer #1

The 5 senior partners has clients x : 6,5,6,5,9

The mean = \sumX/n = 31/5 = 6.2

Mean = 6.2

Now median is the observation which divides the data into two parts

first we have to take data into ascending order

So 5,5,6,6,9

The median = ( no.of observations + 1 )/2 th observation

={(5+1)/2}th​​​​​ observation = 3rd observation

Median = 6

b) The data represents a population.